OpenStudy (anonymous):

Factor completely: 3x^2+17+10 I need some help with the steps to figure this out.

Parth (parthkohli):

= 3x^2 + 27 [Adding 17 + 10] = 3(x^2 + 9) [Factoring]

OpenStudy (anonymous):

ParthKohli, is that the full answer? I'm looking for a complete example so I know how to do the next questions.

Parth (parthkohli):

Yes. that is the answer. You just have to find out the common factors in both and put that out of the bracket. Can you give another so I can explain?

OpenStudy (anonymous):

5y8 - 125

Parth (parthkohli):

Okay. So, a common factor you find in both is 5(125 is divisible by 5)

Parth (parthkohli):

So, you put the 5 out of the bracket. Now what do you have to multiply to 5 to get 5y^8?

Parth (parthkohli):

5 * y^8 = 5y^8, agree? And now 5 * 25 = 125, agree again?

Parth (parthkohli):

So 5(y^8 - 25) is the answer

OpenStudy (anonymous):

Alright, I'm getting the hang of this. Thank you!

Parth (parthkohli):

I can explain more if you wanna perfect this. No problem btw ;)

OpenStudy (anonymous):

There's a2 - 2ab - 15b2 if you don't mind c:

Parth (parthkohli):

Of course :D

Parth (parthkohli):

This is a quadratic expression. It has a different method. Search it on google. Although I'm gonna solve it for you. => a2 -5ab + 3ab - 15b2 => a(a - 5b) + 3b(a - 5b) => (a + 3b)(a - 5b)

Parth (parthkohli):

Do you want me to explain this type of factorization?

OpenStudy (anonymous):

I'll search on google to spare you from typing, haha.

Parth (parthkohli):

Haha....

OpenStudy (anonymous):

How about 4c2 - 12c + 9?

Parth (parthkohli):

Let's see..

Parth (parthkohli):

4c2 - 6c - 6c + 9 = 2c(2c - 3) - 3(2c - 3) = (2c - 3)(2c - 3) \[\LARGE => {(2c - 3)}^{2}\]

Parth (parthkohli):

\[\Huge \text {CAN I HAVE A MEDAL PLEASE?}\]

OpenStudy (anonymous):

How do I give you one?

Parth (parthkohli):

Click that Best Answer thingy in my answer

Parth (parthkohli):

\[\Huge \text {THANKS :D}\]

OpenStudy (anonymous):

No problem c: Thank you!!